OpenStudy (kamille):

Hello. I need some helo There are two vectors: a and b \[a(-\frac{ k \sqrt{3} }{2};\frac{ k }{ 2 };k \sqrt{2})\] \[b(\frac{ k \sqrt{3} }{2};\frac{ k }{ 2 };k \sqrt{2})\] k>0 do i need to find k (according to hat they are lineraly) ?

OpenStudy (kamille):

oh, i forgot the problem: I Need to find the angle between this two vectors!

OpenStudy (ash2326):

Angle between two vectors \(\theta\)= \[\theta =\cos^{-1} (\frac{ a.b}{|a||b|})\] a. b = dot product |a|=magnitude of a |b|=magnitude of b @Kamille Can you try?

OpenStudy (kamille):

what does it mean dot product?

OpenStudy (ash2326):

Suppose you have \[a=2i+3j+4k\] \[b=1i+2j+4k\] \[a.b=2\times3+3\times2+4\times4\]

OpenStudy (kamille):

Oh, okay so \[a.b=(-\frac{ k \sqrt{3}}{ 2 })*(\frac{ k \sqrt{3}}{ 2 })+(\frac{ k }{ 2 })*(\frac{ k }{ 2 })+(k \sqrt{2})(k \sqrt{2})\]

OpenStudy (kamille):

sorry forwaiting, it is still really hard for me to use LATEX :/

OpenStudy (ash2326):

No problem, take your time :) You are doing good

OpenStudy (kamille):

|a|=\[|a|=\sqrt{(-\frac{ k \sqrt{3} }{ 2 })^{2}+(\frac{ k }{ 2 })^{2}+(k \sqrt{2})^{2}}\]

OpenStudy (kamille):

is it right?

OpenStudy (ash2326):

Absolutely :)

OpenStudy (kamille):

thank you for your help! I have understood it now;)

OpenStudy (ash2326):

Welcome :)